OpenStudy (anonymous):

Solve the equation. Check your answer. 4y + 5 = – 31

OpenStudy (anonymous):

where is the other equation

OpenStudy (shaik0124):

4y+5-5=-31+5 4y=-26 y=-26/4

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

that cant be right

OpenStudy (anonymous):

now you do -26/4

OpenStudy (anonymous):

and that will give you what y equals

OpenStudy (shaik0124):

4y+5-5=-31-5 4y=-36 y=-36/4 y=-9

OpenStudy (anonymous):

y=-9

OpenStudy (anonymous):

thanks